# Interpretable Graph Neural Networks for Tabular Data

### v2

@@ 4.2 Evaluation of Explanations
The feature scores produced by IGNNet should ideally reflect the contribution of each feature toward the predicted outcome and, therefore, they should be equivalent to the true Shapley values. As it has been shown that KernelSHAP converges to the true Shapley values when provided with an infinite number of samples (Covert and Lee 2021; Jethani et al. 2022), it is anticipated that the explanations generated by KernelSHAP will progressively converge to more similar values to the scores of IGNNet as the sampling process continues. This convergence arises from KernelSHAP moving towards the true values, while the scores of IGNNet are expected to align with these true values. To examine this conjecture, we explain IGNNet using KernelSHAP and measure the similarity between KernelSHAP’s explanations and IGNNet’s scores following each iteration of data sampling and KernelSHAP evaluation. [^5] KernelSHAP experiments were conducted using the following open-source implementation: https://github.com/iancovert/shapley-regression For the feasibility of the experiment, 500 examples are randomly selected from the test set of each dataset to be explained. The cosine similarity and Spearman rank-order correlation are used to quantify the similarity between explanations. The cosine similarity measures the similarity in the orientation (Han, Kamber, and Pei 2012), while the Spearman rank-order measures the similarity in ranking the importance scores (Rahnama et al. 2021).
@@
The results demonstrate a general trend wherein KernelSHAP’s explanations converge to more similar values to IGNNet’s scores across various data instances and the 35 datasets, as depicted in Figure 3. The consistent convergence to more similar values clearly indicates that IGNNet provides transparent models with feature scores aligned with the true Shapley values. [^6] The complete results of the 35 datasets are provided in the technical appendix.  [^7] We provide illustrations of individual explanations in the technical appendix.
@@ 4.3 Evaluation of Predictive Performance
Detailed results for IGNNet and the five competing algorithms on the 35 datasets are shown in Table 1. The ranking of the five algorithms across 35 datasets, based on their AUC values, reveals OGNNet to exhibit superior performance, claiming the top position, closely followed by IGNNet and XGBoost. In order to investigate whether the observed differences are statistically significant, the Friedman test (Friedman 1939) was employed, which indeed allowed to reject the null hypothesis, i.e., that there is no difference in predictive performance, as measured by the AUC, at the 0.05 level. The result of subsequently applying the post-hoc Nemenyi test (Nemenyi 1963) to determine what pairwise differences are significant, again at the 0.05 level, is summarized in Figure 4. However, the result shows no specific significant pairwise differences between any of the compared algorithms. Furthermore, the results show that using IGNNet instead of the black-box variant, OGNNet, does not significantly reduce the predictive performance while maintaining performance at the level of other powerful algorithms for tabular data, e.g., XGBoost and Random Forests.
@@
4.4 Computational Cost
The computational cost varies based on the architecture of the graph neural network, which can be altered and determined based on the predictive task and the dataset. The cost relative to the conventional graph neural networks remains the same, i.e., the cost does not increase with the proposed approach. The computational cost also depends on the number of features in the dataset. Therefore, the user can decide on the suitable architecture and the acceptable computational cost. Detailed information about the inference time is available in the appendix.
5 Limitations
The proposed approach targets tabular data primarily consisting of numerical features. However, datasets with substantial categorical features impose two challenges. Firstly, Pearson correlation is inadequate in handling correlation between categorical and numerical features or between two categorical features. Secondly, we use one-hot encoding for categorical features, which can result in an exponential growth in dimensionality when dealing with nominal features with numerous categories. Another limitation arises when input features lack correlation, resulting in a completely disconnected graph.
@@ 6 Concluding Remarks
One direction for future research is to explore approaches to model feature interaction in the adjacency matrix that go beyond linear correlations. Understanding how such non-linear interactions between features may impact the model’s interpretability could be an intriguing area of exploration. A second direction is to investigate alternative encoders for categorical features rather than relying on one-hot encoding. It would also be interesting to extend IGNNet to handle non-tabular datasets, including images and text, which would require entirely different approaches to representing each data point as a graph. Another important direction for future work is to use IGNNet for studying possible adversarial attacks on a predictive model. Finally, an important direction would be to complement the empirical evaluation with user-grounded evaluations, e.g., measuring to what extent certain tasks could be more effectively and efficiently solved when the users are provided with a transparent model that shows how the prediction has been computed from the input.
7 Acknowledgement
@@
Appendix A Appendix
Appendix B Illustration of Explanations
In this section, we show how the computed feature scores by IGNNet can be used to understand the feature contributions toward a specific prediction. Note that this is done in exactly the same way as how one would interpret the predictions of a logistic regression model or the feature importance scores generated by the SHAP explainer (Lundberg and Lee 2017). As the computed feature scores reveal exactly how IGNNet formed the prediction, the user can directly see which features have the greatest impact on the final prediction, and possibly also how they may be modified to affect the outcome. To demonstrate this, we present the feature scores for predictions made by IGNNet using two examples from the Adult dataset and the Churn dataset. In the following illustrations, we display the feature scores centered around the bias value, which, when summed with the bias, will produce the exact outcome of IGNNet if the sigmoid function is applied. The scores are sorted according to their absolute values, and only the top 10 features are plotted for ease of presentation. A displayed score $\tau_{i}$ of feature $x_{i}$ represents all the weights and the computations applied to the input value, as shown in equation 5.
$\tau_{i}=$ $\textbf{w}_{i}f(\varphi(\textbf{w}^{(l+1)}(\sum_{u\in\mathcal{N}(i)}\delta_{i,u}\textbf{h}_{u}^{(l)}$ (5)
$+\delta_{i,i}\varphi(\textbf{w}^{(l)}(\sum_{u\in\mathcal{N}(i)}\delta_{i,u}\textbf{h}_{u}^{(l-1)}$
Figure 5: Explanation to a single prediction on Adult dataset.
@@ Appendix B Illustration of Explanations
The first example, from the Adult dataset, shown in Figure 5(a). IGNNet predicted the negative class ($\leq$ 50K) with a narrow margin (0.495). The explanation shows that a single feature (capital-gain=2885) has the highest contribution compared to any other feature value. In the training data, the capital-gain has a maximum value of 99999.0, a minimum value of 0, a mean value of 1068.36, and a 7423.08 standard deviation. To test if the explanation reflects the actual reasoning of IGNNet, we raise the capital-gain value by a smaller value than the standard deviation to be 6000 while leaving the remaining feature values constant, and it turns out to be enough to alter the prediction to a positive ($>$ 50K) with 0.944 as the predicted value. We can also see that the negative score of the capital-gain feature went from -3.67 in the original instance (5(a)) to -0.82 in the modified instance, as shown in Figure 5(b). So the user can adjust the value of an important feature as much as needed to alter the prediction.
The data point, from the Churn dataset, has a positive prediction with a narrow margin (0.565). Consequently, the reduction of the top positively important feature (total day charge), as illustrated in Fig. 6(a), may be enough to obtain a negative prediction. The total day charge has a maximum value of 59.76, a minimum of 0.44, a mean of 30.64, and a 9.17 standard deviation in the training set. However, the total day charge is highly correlated with the total day minutes by more than 0.99, as shown in Fig. 7. The high correlation effect is obvious in the outcome when the total day charge value is reduced by one standard deviation, from 44.56 to 35.39, as the scores of both total day charge and total day minutes drop from 1.06 and 0.99 to 0.72 and 0.5, respectively, as shown in Fig. 6(b). Moreover, the sum of feature scores and bias falls below 0.5 after the sigmoid function, resulting in a negative prediction with a predicted value of 0.297.
Appendix C Ablation Study
@@
Weight on Self-Loop
@@
Threshold on Correlation Values
In the experimental setup, we set a threshold for a correlation value to be added as an edge between two features, which is 0.2 in general and 0.05 in the case of small correlation values. In this set of experiments, we compare the performance of IGNNet with the abovementioned thresholds and without any thresholds. To test the null hypothesis that there is no difference in the predictive performance, as measured by AUC, between the two variations, and since we compare two models, the Wilcoxon signed-rank test (Wilcoxon 1945) is used here. The null hypothesis may be rejected at the 0.05 level, showing that setting a threshold for the edge weight helps with improving the predictive performance of IGNNet. The results are available in Table 2.
The Number of Message-Passing Layers
@@
Appendix D Transparency Evaluation
@@
Appendix E Information about the Used Datasets and Specifications of the Hardware
@@
Appendix F Inference Time
The inference time has been reported using an Intel(R) Core(TM) i9-10885H CPU @ 2.40GHz system with 64.0 GB of RAM, and the GPUs are NVIDIA® GeForce® GTX 1650 Ti with 4 GB GDDR6. The time is reported as the average of 100 different data objects in milliseconds. The results are shown in Table 4.
Table 4: The dataset information.
Dataset Classes Features OHE Nodes Non-Zero Edges Inference Time Abalone 2 8 10 100 3.74 $\pm$ 0.77 Ada Prior 2 14 100 284 4.17 $\pm$ 0.7 Adult 2 14 105 283 4.11 $\pm$ 0.69 Bank 32 nh 2 32 32 38 4.06 $\pm$ 0.72 Covertype 7 54 98 488 9.03 $\pm$ 1.35 Credit Card Fraud 2 30 30 50 3.93 $\pm$ 0.68 Delta Ailerons 2 5 5 11 3.91 $\pm$ 0.67 Electricity 2 8 14 38 3.87 $\pm$ 0.73 Elevators 2 18 18 112 4.31 $\pm$ 0.55 HPC Job Scheduling 4 7 26 258 8.48 $\pm$ 1.81 Fars 8 29 107 2295 8.18 $\pm$ 1.11 1st Order Theorem Prov. 6 51 51 1251 8.34 $\pm$ 0.92 Helena 100 27 27 293 4.61 $\pm$ 0.57 Heloc 2 22 22 370 5.48 $\pm$ 0.51 Higgs 2 28 28 244 6.58 $\pm$ 0.6 Indian Pines 8 220 220 40948 10.7 $\pm$ 1.53 Jannis 4 54 54 542 4.15 $\pm$ 0.7 JM1 2 21 21 415 9.92 $\pm$ 0.79 LHC Identify Jets 5 16 16 184 3.85 $\pm$ 0.76 Madelon 2 500 500 12566 10.42 $\pm$ 2.02 Magic Telescope 2 10 10 54 5.88 $\pm$ 0.53 MC1 2 38 38 1142 4.39 $\pm$ 0.71 Mozilla4 2 5 5 7 4.03 $\pm$ 0.55 Microaggregation2 5 20 20 138 3.82 $\pm$ 0.7 Numerai28.6 2 21 21 231 4.68 $\pm$ 0.618 Otto Group Product 9 93 93 3581 4.06 $\pm$ 0.82 PC2 2 36 36 1040 4.37 $\pm$ 0.73 Phonemes 2 5 5 9 4.31 $\pm$ 0.56 Pollen 2 5 5 15 4.31 $\pm$ 0.67 Satellite 2 36 36 1254 4.84 $\pm$ 0.44 Scene 2 299 304 19178 8.61 $\pm$ 1.27 Speed Dating 2 120 500 3454 7.12 $\pm$ 1.16 Telco Customer Churn 2 19 45 831 5.52 $\pm$ 0.65 Vehicle sensIT 2 100 100 1278 4.66 $\pm$ 0.83 waveform-5000 2 40 40 306 4.62 $\pm$ 0.65

### v3

@@ 4.2 Evaluation of Explanations
The feature scores produced by IGNNet should ideally reflect the contribution of each feature toward the predicted outcome and, therefore, they should be equivalent to the true Shapley values. As it has been shown that KernelSHAP converges to the true Shapley values when provided with an infinite number of samples [8, 22], it is anticipated that the explanations generated by KernelSHAP will progressively converge to more similar values to the scores of IGNNet as the sampling process continues. This convergence arises from KernelSHAP moving towards the true values, while the scores of IGNNet are expected to align with these true values. To examine this conjecture, we explain IGNNet using KernelSHAP and measure the similarity between KernelSHAP’s explanations and IGNNet’s scores following each iteration of data sampling and KernelSHAP evaluation. For the feasibility of the experiment, 500 examples are randomly selected from the test set of each dataset to be explained. The cosine similarity and Spearman rank-order correlation are used to quantify the similarity between explanations. The cosine similarity measures the similarity in the orientation, while the Spearman rank-order measures the similarity in ranking the importance scores [33].
@@
The results demonstrate a general trend wherein KernelSHAP’s explanations converge to more similar values to IGNNet’s scores across various data instances and the 35 datasets, as depicted in Figure 10. The consistent convergence to more similar values clearly indicates that IGNNet provides transparent models with feature scores aligned with the true Shapley values. [^5] The complete results of the 35 datasets are provided in the appendix.
Figure 4: Explanation to a single prediction on Adult dataset.
4.3 Illustration of Explanations
In this section, we show, using a toy example, how the computed feature scores by IGNNet can be used to understand the feature contributions toward a specific prediction. Note that this is done in exactly the same way as how one would interpret the predictions of a logistic regression model or the feature importance scores generated by the SHAP [27] or LIME [34] explainer. As the computed feature scores reveal exactly how IGNNet formed the prediction, the user can directly see which features have the greatest impact on the final prediction, and possibly also how they may be modified to affect the outcome. To demonstrate this, we present the feature scores for predictions made by IGNNet using an example from the Adult dataset. [^6] We provide another example from the Churn dataset in the appendix. In the following illustration, we display the feature scores centered around the bias value, which, when summed with the bias, will produce the exact outcome of IGNNet if the sigmoid function is applied. The scores are sorted according to their absolute values, and only the top 10 features are plotted for ease of presentation. A displayed score $\tau_{i}$ of feature $x_{i}$ represents all the weights and the computations applied to the input value, as shown in equation 5.
$\tau_{i}=$ $w_{i}\mathcal{R}(\varphi(\textbf{w}^{(l)}(\sum_{u\in\mathcal{N}(i)}\delta_{i,u}\textbf{h}_{u}^{(l)}$ (5)
$+\delta_{i,i}\varphi(\textbf{w}^{(l-1)}(\sum_{u\in\mathcal{N}(i)}\delta_{i,u}\textbf{h}_{u}^{(l-1)}$
IGNNet predicted the negative class ($\leq$ 50K) with a narrow margin (0.495). The explanation shows that a single feature (capital-gain=2885) has the highest contribution compared to any other feature value. In the training data, the capital-gain has a maximum value of 99999.0, a minimum value of 0, a mean value of 1068.36, and a 7423.08 standard deviation. Therefore, the capital-gain is relatively low for the selected data instance. To test if the explanation reflects the actual reasoning of IGNNet, we raise the capital-gain value by a smaller value than the standard deviation to be 6000 while leaving the remaining feature values constant, and it turns out to be enough to alter the prediction to a positive ($>$ 50K) with 0.944 as the predicted value. We can also see that the negative score of the capital-gain feature went from -3.67 in the original instance (4(a)) to -0.82 in the modified instance, as shown in Figure 4(b). So the user can adjust the value of an important feature as much as needed to alter the prediction.
@@ 4.4 Evaluation of Predictive Performance
Detailed results for IGNNet and the five competing algorithms on the 35 datasets are shown in Table 1. The ranking of the six algorithms across 35 datasets, based on their AUC values, reveals OGNNet to exhibit superior performance, claiming the top position, closely followed by IGNNet and XGBoost. In order to investigate whether the observed differences are statistically significant, the Friedman test [16] was employed, which indeed allowed to reject the null hypothesis, i.e., reject that there is no difference in predictive performance, as measured by the AUC, at the 0.05 level. The result of subsequently applying the post-hoc Nemenyi test [30] to determine what pairwise differences are significant, again at the 0.05 level, is summarized in Figure 5. However, the result shows no specific significant pairwise differences between any of the compared algorithms. Furthermore, the results show that using IGNNet instead of the black-box variant, OGNNet, does not significantly reduce the predictive performance while maintaining performance at the level of other powerful algorithms for tabular data, e.g., XGBoost and Random Forests.
4.5 Computational Cost
The computational cost varies based on the architecture of the graph neural network, which can be altered and determined based on the predictive task and the dataset. The cost relative to the conventional graph neural networks remains the same, i.e., the cost does not increase with the proposed approach. The computational cost also depends on the number of features in the dataset. Therefore, the user can decide on the suitable architecture and the acceptable computational cost.
5 Limitations
The proposed approach targets tabular data primarily consisting of numerical features. However, datasets with substantial categorical features impose two challenges. Firstly, Pearson correlation is inadequate in handling correlation between categorical and numerical features or between two categorical features. Secondly, we use one-hot encoding for categorical features, which can result in an exponential growth in dimensionality when dealing with nominal features with numerous categories. Another limitation arises when input features lack correlation, resulting in a null graph (a completely disconnected graph).
@@ 6 Concluding Remarks
One direction for future research is to explore approaches to model feature interactions in the adjacency matrix that go beyond linear correlations. Understanding how such non-linear interactions between features may impact the model’s interpretability could be an intriguing area of exploration. A second direction is to investigate alternative encoders for categorical features rather than relying on one-hot encoding. It would also be interesting to extend IGNNet to handle non-tabular datasets, including images and text, which would require entirely different approaches to representing each data point as a graph. Another important direction for future work is to use IGNNet for studying possible adversarial attacks on a predictive model. Finally, an important direction would be to complement the empirical evaluation with user-grounded evaluations, e.g., measuring to what extent certain tasks could be more effectively and efficiently solved when the users are provided with a transparent model that shows how the prediction has been computed from the input.
{ack}
@@
Appendix A Illustration of Explanations
@@
In this section, we illustrate the computed feature scores by IGNNet on one example from the Churn dataset and show how the explanation can reflect the interaction between features. The data point, from the Churn dataset, has a positive prediction with a narrow margin (0.565). Consequently, the reduction of the top positively important feature (total day charge), as illustrated in Fig. 6(a), may be enough to obtain a negative prediction. The total day charge has a maximum value of 59.76, a minimum of 0.44, a mean of 30.64, and a 9.17 standard deviation in the training set. However, the total day charge is highly correlated with the total day minutes by more than 0.99, as shown in Fig. 7. The high correlation effect is obvious in the outcome when the total day charge value is reduced by one standard deviation, from 44.56 to 35.39, as the scores of both total day charge and total day minutes drop from 1.06 and 0.99 to 0.72 and 0.5, respectively, as shown in Fig. 6(b). Moreover, the sum of feature scores and bias falls below 0.5 after the sigmoid function, resulting in a negative prediction with a predicted value of 0.297.
@@
Appendix B Ablation Study
@@
B.1 Weight on Self-Loop
@@
B.2 Threshold on Correlation Values
In the experimental setup, we set a threshold for a correlation value to be added as an edge between two features, which is 0.2 in general and 0.05 in the case of small correlation values. In this set of experiments, we compare the performance of IGNNet with the abovementioned thresholds and without any thresholds. To test the null hypothesis that there is no difference in the predictive performance, as measured by AUC, between the two variations, and since we compare two models, the Wilcoxon signed-rank test [39] is used here. The null hypothesis may be rejected at the 0.05 level, showing that setting a threshold for the edge weight helps with improving the predictive performance of IGNNet. The results are available in Table 2.
B.3 The Number of Message-Passing Layers
@@
Appendix C Transparency Evaluation
@@
Appendix D Information about the Used Datasets and Specifications of the Hardware
@@
Appendix E Inference Time
The inference time has been reported using an Intel(R) Core(TM) i9-10885H CPU @ 2.40GHz system with 64.0 GB of RAM, and the GPU is NVIDIA® GeForce® GTX 1650 Ti with 4 GB GDDR6. The time is reported as the average over 100 different data objects in milliseconds. The results are shown in Table 4. Moreover, Table 4 shows the original number of features (Features), the subsequent number of nodes after the binarization of categorical features using one-hot-encoding (OHE Nodes), and the number of non-zero edges in the graph.
Table 4: The inference time across different datasets.
Dataset Classes Features OHE Nodes Non-Zero Edges Inference Time (ms) Abalone 2 8 10 100 3.74 $\pm$ 0.77 Ada Prior 2 14 100 284 4.17 $\pm$ 0.7 Adult 2 14 105 283 4.11 $\pm$ 0.69 Bank 32 nh 2 32 32 38 4.06 $\pm$ 0.72 Covertype 7 54 98 488 9.03 $\pm$ 1.35 Credit Card Fraud 2 30 30 50 3.93 $\pm$ 0.68 Delta Ailerons 2 5 5 11 3.91 $\pm$ 0.67 Electricity 2 8 14 38 3.87 $\pm$ 0.73 Elevators 2 18 18 112 4.31 $\pm$ 0.55 HPC Job Scheduling 4 7 26 258 8.48 $\pm$ 1.81 Fars 8 29 107 2295 8.18 $\pm$ 1.11 1st Order Theorem Prov. 6 51 51 1251 8.34 $\pm$ 0.92 Helena 100 27 27 293 4.61 $\pm$ 0.57 Heloc 2 22 22 370 5.48 $\pm$ 0.51 Higgs 2 28 28 244 6.58 $\pm$ 0.6 Indian Pines 8 220 220 40948 10.7 $\pm$ 1.53 Jannis 4 54 54 542 4.15 $\pm$ 0.7 JM1 2 21 21 415 9.92 $\pm$ 0.79 LHC Identify Jets 5 16 16 184 3.85 $\pm$ 0.76 Madelon 2 500 500 12566 10.42 $\pm$ 2.02 Magic Telescope 2 10 10 54 5.88 $\pm$ 0.53 MC1 2 38 38 1142 4.39 $\pm$ 0.71 Mozilla4 2 5 5 7 4.03 $\pm$ 0.55 Microaggregation2 5 20 20 138 3.82 $\pm$ 0.7 Numerai28.6 2 21 21 231 4.68 $\pm$ 0.618 Otto Group Product 9 93 93 3581 4.06 $\pm$ 0.82 PC2 2 36 36 1040 4.37 $\pm$ 0.73 Phonemes 2 5 5 9 4.31 $\pm$ 0.56 Pollen 2 5 5 15 4.31 $\pm$ 0.67 Satellite 2 36 36 1254 4.84 $\pm$ 0.44 Scene 2 299 304 19178 8.61 $\pm$ 1.27 Speed Dating 2 120 500 3454 7.12 $\pm$ 1.16 Telco Customer Churn 2 19 45 831 5.52 $\pm$ 0.65 Vehicle sensIT 2 100 100 1278 4.66 $\pm$ 0.83 waveform-5000 2 40 40 306 4.62 $\pm$ 0.65